A solution of table sugar in water contains 5.25% by mass sucrose. Calculate:



1) the mass of sucrose in 225 g of solution



2) the amount of solution needed to obtain 70.0 g of sucrose



3)the mass of water contained in 450 g of the solution



4)the mole fraction of sucrose and water in 5.25% by mass sucrose.

1
Expert's answer
2022-03-16T14:25:03-0400

(1) 100 g sucrose solution has 5.25 g of sucrose

so, 225 g of solution has mass of sucrose = "\frac{5.25}{100}\times 225 = 11.8125 g"


(2) 5.25 g of sucrose present in 100 g solution

so, 70 g of sucrose present in "\frac{100}{5.25}\times 70 = 1333.33 g"


(3) 100 g of solution has 5.25 g sucrose

450 g of solution has = 21 g sucrose

Mass of water = 429 g


(4)mass of water in 5.25 % = 94.75 g

mole of water = "\frac{94.75}{18}=5.264"

mole of sucrose = "\frac{5.25}{344}=0.015"

Mole fraction o sucrose

"\chi_1= \frac{n_1}{n_1+n_2}= \frac{0.015}{0.015+5.264}=0.0028"


Mole fraction of water

"\chi_2= 1-0.0028 =0.9971"